# Stabilizer Entanglement as a Magic Highway

### v2

@@ Abstract
Non-stabilizerness is a key resource for fault-tolerant quantum computation, yet its interplay with entanglement in dynamical settings remains underexplored. We study a well-controlled, analytically tractable setup that isolates entanglement generation from magic injection. We analytically and numerically demonstrate that stabilizer entanglement functions as a highway that facilitates the spreading of locally injected magic throughout the entire system. Specifically, for an initial stabilizer state with bipartite entanglement $E$, the total magic growth, quantified by the linear stabilizer entropy $Y$, follows $\overline{Y}\propto 2^{-|A|-E}$ under a Haar random unitary on a local subregion $A$. Moreover, when applying a tensor product of local Haar random unitaries, the resulting state’s global magic approaches that of a genuine Haar random state if the initial stabilizer state is sufficiently entangled by a system-size-independent amount. Similar results are also obtained for tripartite stabilizer entanglement. We further extend our analysis to non-stabilizer entanglement and magic injection via a shallow-depth brickwork circuit, and find that the qualitative picture of our conclusion remains unchanged.
@@
Introduction.- Understanding what distinguishes quantum systems from classical ones has been a challenging and multi-faceted endeavor. Much of the impact of quantumness in physics has been focusing on the idea of quantum entanglement, which played a central role in characterizing topological order [1, 2, 3], dynamically induced phases [4], quantum chaos, and even the emergence of spacetime [5]. However, entanglement being a quantum correlation, is just one facet of quantumness. Quantum advantage, for example, is concerned with the hardness of simulating quantum systems on a classical computer. This facet is distinct from entanglement as many highly entangled systems prepared using Clifford operations violate Bell’s inequalities but can be efficiently simulated classically [6, 7]. The notion of classical hardness therefore constitutes a second layer of quantumness and is intimately connected to non-stabilizerness or magic [8, 9, 10, 11, 12, 13].
@@
Like entanglement, magic is an important but distinct quantum resource for realizing fault-tolerant universal quantum computation [13, 14]. Therefore, it is crucial to understand any physical consequences played by this orthogonal dimension to entanglement in quantum systems. The importance of non-stabilizerness has been largely overlooked in quantum many-body physics until recently, where a flurry of activities have been dedicated to study its total extent in various systems, which has implication for resource estimation in state preparation and quantum simulation [15, 16, 17, 18, 19, 20], in characterizing phase transitions [21, 22], in quantum dynamics [23, 24, 25, 26, 27], in improved quantum simulation with tensor networks [28, 29], and in quantum gravity [30, 31]. Arguably, the most fascinating perspective comes from the interplay between magic and entanglement [32, 31, 33, 34, 35, 36, 27] as both are needed for quantum advantage.
Recent work revealed that, unlike the spread of entanglement which takes time linear in system size in spatially local Haar random circuits, magic appears to saturate in $O(\log n)$ time [37] (whereas it spreads much more slowly in Hamiltonian systems [38]). In the long time limit, it is also understood that magic in these highly entangled system is similar to the scrambling/encoding of information in quantum error correcting codes where it quickly delocalizes and can only be found when one can access a large part of the system [39, 26, 40]. This asymptotic behavior is consistent with the observation that subsystem magic quickly rises and decays to almost zero as a state thermalizes [24].
Despite these preliminary findings, the fundamental mechanisms underlying magic dynamics, particularly the interplay between magic injection and entanglement, remain poorly understood. Specifically, the role of entanglement in facilitating the growth and spread of magic, and vice versa, is yet to be determined in a quantitative manner. In generic quantum circuits and Hamiltonian dynamics, this interplay is inherently complex, as each time step of evolution simultaneously injects magic and increases entanglement [41, 37]. To gain deeper insights into magic dynamics, a well-controlled setup that isolates different contributing factors is essential. Here we study precisely one such setting and show that entanglement, or at least stabilizer entanglement, acts as a conduit for magic, allowing it to spread much faster to parts of the system that are entangled. This would constitute an important first step in understanding how magic dynamics is distinct and yet interlinked with various types of entanglement dynamics.
In this work, we analytically and numerically demonstrate that, given an initial stabilizer state with bipartite entanglement $E$, the total growth of magic, as quantified by the linear stabilizer entropy $Y$ under magic injection via a Haar-random unitary acting on a local subregion $A$, is enhanced by the presence of entanglement: $\overline{Y}\propto 2^{-|A|-E}$. Furthermore, when a unitary of the form $U=U_{A}\otimes U_{B}$, consisting of independently Haar-random local unitaries on subsystems $A$ and $B$, is applied, the global magic of the resulting state becomes nearly indistinguishable from that of a fully Haar-random state, provided the initial state is sufficiently entangled. Interestingly, the amount of entanglement necessary for such indistinguishability from Haar random magic is independent of system size. Similar results are also obtained for tripartite stabilizer entanglement. We further extend our analysis to non-stabilizer entanglement and magic injection via a shallow-depth brickwork circuit, and find that the qualitative picture of our conclusion remains unchanged in these cases.
@@
Quantum magic and its measure.- Quantum magic quantifies the amount of “non-stabilizer” or “non-Clifford” resources in a quantum state. A computable measure for quantum magic, which we will primarily focus on in this work, is given by the linear stabilizer entropy. Denote by $P_{N}$ the set of $N$-qubit Pauli strings. The linear stabilizer entropy for a pure state $\rho=|\psi\rangle\langle\psi|$ is defined as $Y(\rho):=2^{-N}\sum_{P\in P_{N}}{\rm tr}(P\rho)^{4}$. To gain some intuition about this quantity, first consider a stabilizer pure state. There are precisely $2^{N}$ Pauli strings corresponding to the stabilizer group elements of $\rho$ for which ${\rm tr}(P\rho)^{4}=1$, and ${\rm tr}(P\rho)^{4}=0$ for all other Pauli strings. Thus $Y(\rho)=1$ for stabilizer states. For a generic non-stabilizer pure state, the Pauli spectrum satisfies the normalization condition: $12^{-N}\sum_{P\in P_{N}}{\rm tr}(P\rho)^{2}=$ 1. Therefore, $Y(\rho)$ defined above can be thought of as the second moment of the Pauli spectrum of $\rho$, which characterizes how delocalized the expectation values of $P$ are over all possible Pauli strings. The linear stabilizer entropy satisfies the following properties: (i) faithfulness: let $\text{STAB}_{n}$ denote the set of all $n$-qubit pure stabilizer states, then $Y(\rho)=1$ iff $\rho\in\text{STAB}_{n}$, otherwise $Y(\rho)<1$; (ii) invariant under Clifford unitaries $U_{C}$: $Y(U_{C}\rho U^{\dagger}_{C})=Y(\rho)$; (iii) multiplicative: $Y(\rho_{A}\otimes\rho_{B})=Y(\rho_{A})Y(\rho_{B})$; (iv) the closely related quantity $M^{\rm lin}:=1-Y$ is a strong pure-state monotone under free operations [42]. Hence, $Y(\rho)$ serves as a good measure for quantum magic from the point of view of magic-state resource theory. Moreover, the logarithm of $Y(\rho)$ is related to the second stabilizer Rényi entropy $M_{2}(\rho)=-{\rm log}Y(\rho)$, another widely used magic measure [43].
@@
Figure 1: Schematics of the setup used for analytical and numerical calculations. The initial states are bipartite or tripartite stabilizer pure states. (a) A Haar random unitary injects magic by acting on a subregion $A$. (b) A factorized unitary $U=U_{A}\otimes U_{B}$, where $U_{A}$ and $U_{B}$ are independently Haar-random, acts on the initial stabilizer state. (c) A factorized unitary $U=U_{A}\otimes U_{B}$ injects magic by acting on subregion $AB$ of a tripartite system. (d) Initial state (2) for numerical simulations of the setup in (a). The state consists of $f_{A}$ and $f_{B}$ qubits in |+ket|$\rangle$| + ⟩ state in subsystem $A$ and $B$, respectively, along with $E$ Bell pairs shared between $A$ and $B$. (e) Initial state for obtaining Eq. (5). The state initial state consists of $g$ GHZ states shared among $A$, $B$, and $C$; $b_{AB}$, $b_{AC}$, and $b_{BC}$ Bell pairs shared between subsystems $AB$, $AC$, and $BC$, respectively, and $f_{A}$, $f_{B}$, $f_{C}$ single-qubit states in each individual subsystem.
Stabilizer entanglement facilitates magic injection.-To illustrate how stabilizer entanglement can influence magic injection, we examine the effect of a single-qubit $T$ gate on two distinct stabilizer states: an unentangled product state, $|\psi_{1}\rangle=|+\rangle^{\otimes N}=\left((|0\rangle+|1\rangle)/\sqrt{2}\right)^{\otimes N}$, and a GHZ state with long-range entanglement, $|\psi_{2}\rangle=(|0\rangle^{\otimes N}+|1\rangle^{\otimes N})/\sqrt{2}$. In both cases, the initial states contain no magic, and the $T$ gate injects magic locally. However, the manner in which this injected magic manifests differs significantly between the two cases. In $T|\psi_{1}\rangle$, the injected magic remains locally accessible and can be directly probed at the specific qubit where the $T$ gate is applied. In contrast, for $T|\psi_{2}\rangle$, the state appears non-magical to any local observer with access only to a subregion of the system. This simple example illustrates that the propagation of locally injected magic throughout the system is intrinsically tied to the entanglement structure of the initial state.
To further quantify how stabilizer entanglement facilitates magic spreading, we consider the setup illustrated in Fig. 1(a), which enables an analytical calculation. Specifically, we begin with an initial stabilizer pure state $\rho=|\psi\rangle\langle\psi|$, partitioned into subregion $A$ and its complement $B$. A Haar random unitary $U_{A}$ is then applied to subregion $A$, and we are interested in computing the linear stabilizer entropy of the resulting state $\sigma=U_{A}|\psi\rangle\langle\psi|U_{A}^{\dagger}$, averaged over Haar random unitaries. We first make a few remarks on this particular choice of setup before showing the result. This construction is a natural generalization of the toy example discussed above, replacing the single-qubit unitary with a Haar random unitary on a finite subregion. While the Haar-random unitary may appear nonlocal within subregion $A$, it is local relative to the entire system if $|A|\ll|B|$. The use of Haar random unitaries primarily serves to render the analytical calculations more tractable. We will show below that key qualitative features of our results persist even when considering more structured unitaries within subregion $A$. Secondly, this setup clearly separates entanglement generation with respect to the bipartition and magic injection. The unitary acting on $A$ injects magic to the system while preserving the amount of entanglement between $A$ and $B$, including the full entanglement spectrum. As many, if not most, operationally convenient circuits apply Clifford gates and non-Clifford phase gates in separate stages, e.g. Clifford+T decomposition [13] and the PFC ensemble that prepare pseudorandom unitaries [44], our setup is to probe the internal dynamics that take place between the most elementary layers in those circuits. This is complementary to the random unitary circuit model used in Ref. [37] to study magic saturation. In that model, unitary gates simultaneously generate entanglement and inject magic, leading to universal late-time magic dynamics prior to saturation, regardless of the initial state.
@@
We analytically compute the linear stabilizer entropy $Y(U_{A}|\psi\rangle\langle\psi|U_{A}^{\dagger})$, averaged over the ensemble of Haar random unitaries $U_{A}$. In the limit $|A|\gg 1$, we obtain, to leading order,
$\overline{Y}=\mathbb{E}_{U_{A}}\ Y\left(U_{A}|\psi\rangle\langle\psi|\ U_{A}^{\dagger}\right)=4\cdot{{2}^{-|A|-E}}\left[1+O({2}^{-{|A|-E}})\right],$ (1)
where $E=-\rm{tr}(\rho_{A}\text{log}_{2}\rho_{A})$ is the amount of entanglement between $A$ and $B$, which is an integer for a stabilizer state $|\psi\rangle$. An exact calculation of $\overline{Y}$ for arbitrary $|A|$ and $E$ is possible, and is shown in the Supplemental Material (SM) [45]. Eq. (1) establishes a quantitative relation between the efficiency of local magic injection and the amount of stabilizer entanglement contained in the target state. For a fixed subregion $A$, the average $Y(\rho)$ decays exponentially with the amount of entanglement between $A$ and its complement. Therefore, global magic injection on $AB$ is enhanced with entanglement, which acts as a throttle that controls how much magic can flow to region $B$ from $A$. A detailed derivation of Eq. (1) is provided in the SM [45], yet its qualitative features, particularly the factor $2^{-|A|-E}$ can be understood straightforwardly. Since the unitary acts only on subsystem $A$, stabilizers supported entirely on $B$ remains unchanged under $U_{A}$ and contribute to $Y(\rho)$. A simple counting shows that the number of surviving stabilizers is precisely $2^{|B|-E}$, leading to the factor $2^{-|A|-E}$ in Eq. (1). The prefactor 4 in front of $2^{-|A|-E}$ is, in some sense, non-universal. To see this, let us define the generalized linear stabilizer entropy $Y_{\alpha}(\rho):=2^{-N}\sum_{P\in P_{N}}{\rm tr}(P\rho)^{2\alpha}$. Eq. (1) thus corresponds to the specific case $\alpha=2$. Similarly, $Y_{\alpha}(\rho)$ is related to the $\alpha$-th stabilizer Rényi entropy: $M_{\alpha}(\rho)=\frac{1}{1-\alpha}{\rm log}Y_{\alpha}(\rho)$. In the limit $\alpha\rightarrow\infty$, only the surviving stabilizers supported on $B$ contribute, yielding $\overline{Y_{\infty}}=2^{-|A|-E}$. For arbitrary $\alpha$, we have $\overline{Y_{\alpha}}=[1+c_{\alpha}(|A|,E)]2^{-|A|-E}$, with the one coming from the surviving stabilizers, and $c_{\alpha}(|A|,E)$ coming from contributions of all other Pauli strings, which in general depends on both $|A|$ and $E$ [46]. Specifically, Eq. (1) implies that for $\alpha=2$, we have $c_{2}=3$.
Figure 2: (a) Numerical simulations of the setup depicted in Fig. 1(a). We consider initial states of the form Eq. (2) with varying $|A|$ and $E$. We compute the average value of $\overline{Y}$ (upper triangles), as well as $2^{-\overline{M_{2}}}$ (lower triangles), where $\overline{M_{2}}$ is the average value of the second stabilizer Rényi entropy. The analytical result Eq. (1) is shown as dashed lines. (b) Numerical simulations of the setup depicted in Fig. 1(b). Choice of the initial state is the same as in (a). The analytic result Eq. (LABEL:eq:UaUb) is shown as dashed lines. (c) Exact results (squares) and leading order contributions Eq. (5) of the setup depicted in Fig. 1(c). (d) Same as in (a) but for non-stabilizer initial states with $k$ imperfect Bell pairs $|\phi_{\theta}\rangle=\cos\theta|00\rangle+\sin\theta|11\rangle$ shared between $A$ and $B$. We compute the magic already existing in the initial state $\overline{Y}_{\text{ini}}$ (dashed lines) and the change after the action of $U_{A}$: $\overline{Y}_{\text{final}}/\overline{Y}_{\text{ini}}$ (solid lines). (e) Same as in (c) but initial states are now constructed with four different types of entanglement spectra (see main text). Error bars are smaller than the size of the data points and hence not visible. (f) Same as in (d)&(e), but now plotting $\overline{Y}_{\rm final}$ in each case considered.
To test Eq. (1), particularly the regime of validity of neglecting subleading corrections, we perform numerical simulations of the setup depicted in Fig. 1(a). Recall that any stabilizer state $|\psi\rangle$ can be brought to the form of Eq. (2) by local Clifford unitaries [47]: $|\Phi\rangle=U^{\rm Cliff}_{A}\otimes U^{\rm Cliff}_{B}|\psi\rangle$ such that
$|\Phi\rangle=|+\rangle^{\otimes f_{A}}\ |{\rm Bell}\rangle_{AB}^{\otimes E}\ |+\rangle^{\otimes f_{B}},$ (2)
which is illustrated in Fig. 1(d). The state consists of $f_{A}$ and $f_{B}$ qubits in |+ket|$\rangle$| + ⟩ state in subsystem $A$ and $B$, respectively, along with $E$ Bell pairs shared between $A$ and $B$, so that the entanglement entropy between $A$ and $B$ is $E$. Since $\overline{Y}$ is obtained by averaging over the Haar random ensemble, $U^{\rm Cliff}_{A}$ can be simply absorbed into the average. Moreover, $Y(\rho)$ is invariant under Clifford unitary $U_{B}^{\rm Cliff}$. Thus we have
$\mathbb{E}_{U_{A}}Y\left(U_{A}|\psi\rangle\langle\psi|U_{A}^{\dagger}\right)=\mathbb{E}_{U_{A}}Y\left(U_{A}|\Phi\rangle\langle\Phi|U_{A}^{\dagger}\right)$ (3)
for an arbitrary stabilizer state $|\psi\rangle$. We numerically average over $Y(U_{A}|\Phi\rangle\langle\Phi|U_{A}^{\dagger})$ with Haar random $U_{A}$ for various values of $|A|$ and $E$. The results are shown in Fig. 2(a), where we find excellent agreement with the analytical prediction of Eq. (1). In particular, the subleading corrections become negligible already at moderate subsystem sizes and entanglement $|A|+E\gtrsim 4$. We also simulate the Haar averaged second stabilizer Rényi entropy $\overline{M_{2}}$, which turns out to be extremely close to ${\rm log}\ {\overline{Y}}$. This implies that the average in Eq. (1) is also typical, as sample-to-sample fluctuations are extremely small. The same qualitative behavior holds if $U_{A}$ is given by a brickwork circuit instead of Haar random unitary, but the total injected magic is limited by the depth (see SM [45]).
@@
One can further explore the most general approach to magic injection while preserving the entanglement spectrum. Starting from a stabilizer state, we now apply a unitary that is a tensor product of local unitaries acting on subsytems $A$ and $B$: $U=U_{A}\otimes U_{B}$, as illustrated in Fig. 1(b). We compute the linear stabilizer entropy of the resulting state, averaged over Haar random $U_{A}$ and $U_{B}$. In the limit where $|A|\gg 1$ and $|B|\gg 1$, we find [45]
$\overline{Y}$ $=$ $\mathbb{E}_{U_{A}}\mathbb{E}_{U_{B}}Y\left(U_{A}U_{B}|\psi\rangle\langle\psi|U_{A}^{\dagger}U_{B}^{\dagger}\right)$
$=$ $4\cdot{2}^{-N}\left[1+3\cdot 2^{-2E}+O({{2}^{-|A|-E}})+O({{2}^{-|B|-E}})\right],$

### v3

# Stabilizer Entanglement Enhances Magic Injection

@@ Abstract
Non-stabilizerness is a key resource for fault-tolerant quantum computation, yet its interplay with entanglement in dynamical settings remains underexplored. We address this by analyzing a well-controlled, analytically tractable setup, where we show that entanglement acts as a conduit that teleports magic across the system, thereby enhancing magic injection. Using exact calculations, we prove that when a Haar-random unitary $U_{A}$ is applied to a subsystem $A$ of an entangled stabilizer state, the total injected magic increases with the entanglement between $A$ and its complement. More generally, for any unitary $U_{A}$, we show that this enhancement is maximized when $A$ is maximally entangled with its complement, in which case the total injected magic is exactly given by the unitary stabilizer Rényi entropy we introduce. This quantity provides both a directly computable measure of unitary magic and a lower bound on the minimum number of $T$ gates required to synthesize $U_{A}$. We further extend our analysis to tripartite stabilizer entanglement, non-stabilizer entanglement, and magic injection via shallow-depth brickwork circuits, finding that the qualitative picture remains unchanged.
@@
Introduction.- Understanding what distinguishes quantum systems from classical ones has been a challenging and multi-faceted endeavor. Much of the study of quantumness in physics has focused on the idea of quantum entanglement, which played a central role in characterizing topological order [1, 2, 3], dynamically induced phases [4], quantum chaos, and even the emergence of spacetime [5]. However, entanglement being a quantum correlation, is just one facet of quantumness. Quantum advantage, for example, is concerned with the hardness of simulating quantum systems on a classical computer. This facet is not captured by entanglement alone as many highly entangled systems prepared using Clifford operations violate Bell’s inequalities but can be efficiently simulated classically [6, 7]. The notion of classical hardness therefore constitutes a second layer of quantumness and is intimately connected to non-stabilizerness or magic [8, 9, 10, 11, 12, 13].
@@
Like entanglement, magic is an important but distinct quantum resource for realizing fault-tolerant universal quantum computation [13, 14]. The importance of non-stabilizerness has been largely overlooked in quantum many-body physics until recently, where a flurry of activities have been dedicated to study its total extent in various systems, which has implication for resource estimation in state preparation and quantum simulation [15, 16, 17, 18, 19, 20, 21, 22], in characterizing phase transitions [23, 24], in quantum dynamics [25, 26, 27, 28, 29, 30], in improved quantum simulation with tensor networks [31, 32, 33, 34, 35], and in quantum gravity [36, 37]. Arguably, the most fascinating perspective comes from the interplay between magic and entanglement [38, 37, 39, 40, 41, 42, 29] as both are needed for quantum advantage. Recent work revealed that magic in highly entangled systems is similar to the scrambling/encoding of information in quantum error correcting codes where it quickly delocalizes and can only be found when one can access a large part of the system [43, 28, 44]. This asymptotic behavior is consistent with the observation that subsystem magic quickly rises and decays to almost zero as the system thermalizes [26].
Despite recent progress, the fundamental mechanisms governing magic dynamics, particularly the interplay between magic injection and entanglement, remain poorly understood. In generic quantum circuits and Hamiltonian dynamics, entanglement and magic dynamics are interwined, as each time step of evolution simultaneously injects magic and increases entanglement [45, 46]. To disentangle their roles, it is essential to study a controlled setting where one can isolate their contributions. In this work, by constructing such a setup, we aim to address the following question: how does entanglement affect the injection of magic into a quantum system?
We show that preexisting stabilizer entanglement (i.e. entanglement of a nonmagical, stabilizer state) can enhance magic injection via gate teleportation. More explicitly, using exact calculations, we prove that when a Haar-random unitary $U_{A}$ is applied to a subsystem $A$ of an entangled stabilizer state, the total injected magic increases with the entanglement between $A$ and its complement. More generally, for any unitary $U_{A}$, we show that this enhancement is maximized when $A$ is maximally entangled with its complement, in which case the total injected magic is exactly given by the unitary stabilizer Rényi entropy we introduce. This quantity provides both a directly computable measure of unitary magic and a lower bound on the minimum number of $T$ gates (defined as $T=e^{i\frac{\pi}{8}Z}$) required to synthesize $U_{A}$, also known as $T$-count. We further extend our analysis to tripartite stabilizer entanglement, non-stabilizer entanglement, and magic injection via shallow-depth brickwork circuits, finding that the qualitative picture remains unchanged.
Quantum magic and its measure.- Clifford operations consist of unitaries that map Pauli operators to Pauli operators, the preparation of Pauli eigenstates, and the measurement in the Pauli basis among other things. Their simplicity around Pauli operators allows them to be efficiently simulable on a classical computer. The states that can be prepared with Clifford operations are called stabilizer states. A stabilizer pure state $|\psi\rangle$ is one where $s|\psi\rangle=|\psi\rangle$ for all $s\in S$ where $S$ is a size-$2^{N}$ Abelian subgroup of the Pauli group $P_{N}$ over $N$ qubits. It can be shown that $\rho=|\psi\rangle\langle\psi|=\frac{1}{2^{N}}\sum_{s\in S}s$ — a stabilizer pure state is an equal superposition of the $2^{N}$ stabilizer group elements.
Quantum magic quantifies the amount of “non-Clifford” or “non-stabilizer” resources needed in a quantum state or process. In this work, we focus on a computable measure of magic in quantum states called linear stabilizer entropy [47] defined as
$Y^{\rm lin}(\rho):=1-\frac{1}{2^{N}}\sum_{P\in P_{N}}{\rm tr}(P\rho)^{4}.$ (1)
[…]
@@
The linear stabilizer entropy satisfies the following properties: (i) faithfulness: let $\text{STAB}_{N}$ denote the set of all $N$-qubit pure stabilizer states, then $Y^{\rm lin}(\rho)=0$ iff $\rho\in\text{STAB}_{N}$, otherwise $Y^{\rm lin}(\rho)>0$; (ii) invariant under Clifford unitaries $U_{C}$: $Y^{\rm lin}(U_{C}\rho U^{\dagger}_{C})=Y^{\rm lin}(\rho)$; (iii) monotonicity: $Y^{\rm lin}$ is a strong pure-state monotone under free operations [48]. Hence, $Y^{\rm lin}(\rho)$ serves as a good measure for quantum magic from the point of view of magic-state resource theory, which increases and approaches unity as magic increases. Notice that $Y^{\rm lin}(\rho)$ is also closely related to the second stabilizer Rényi entropy (SRE) $M_{2}(\rho)=-{\rm log}(1-Y^{\rm lin}(\rho))$, another widely used magic measure [47].
@@
Figure 1: Schematics of the setup considered in this work. The initial states are bipartite stabilizer pure states. (a) A Haar random unitary injects magic by acting on a subregion $A$. (b) A factorized unitary $U=U_{A}\otimes U_{B}$, where $U_{A}$ and $U_{B}$ are independently Haar-random, acts on the initial stabilizer state. (c) Schematic of gate teleportation. Because of the Bell pairs between $A$ and $B$, $U_{A}$ acting on $A$ is equivalent to $U^{\prime}_{A}$} and $U^{\prime}_{B}$} acting on $A$ and $B$.
Stabilizer entanglement enhances magic injection.- To see how entanglement impacts magic injection, consider the setup illustrated in Fig. 1(a). We take an initial stabilizer pure state $\rho$, partitioned into subregion $A$ and its complement $B$. A Haar random unitary $U_{A}$ is then applied to subregion $A$, and we are interested in computing the linear stabilizer entropy of the resulting state $\sigma=U_{A}|\psi\rangle\langle\psi|U_{A}^{\dagger}$, averaged over Haar random unitaries. This setup preserves the entanglement spectrum between $A$ and $B$ and models the process where magic is locally injected by a non-Clifford unitary gate to a stabilizer state when $|A|\ll|B|$. We will show in SM [49] that key qualitative features of our results also persist even when considering more structured unitaries within subregion $A$.
@@
We then compute Haar-averaged the linear stabilizer entropy $Y^{\rm lin}(U_{A}|\psi\rangle\langle\psi|U_{A}^{\dagger})$ exactly. To leading order,
$\overline{Y^{\rm lin}}$ $=$ $\mathbb{E}_{U_{A}}\ Y^{\rm lin}\left(U_{A}|\psi\rangle\langle\psi|\ U_{A}^{\dagger}\right)$ (2)
$=$ $1-4\cdot{{2}^{-|A|-E}}\left[1+O({2}^{-{|A|-E}})\right],$
where $|A|\gg 1$ and $E=-\rm{tr}(\rho_{A}\text{log}\rho_{A})$ is the amount of entanglement between $A$ and $B$, which is an integer for a stabilizer state. An exact closed-form expression for $\overline{Y^{\rm lin}}$ at arbitrary $|A|$ and $E$ is derived in the Supplemental Material (SM) [49]. It turns out that the leading-order analytical formula already provides an excellent approximation for moderate system sizes and entanglement: numerical simulations show deviations become negligible once $|A|+E\gtrsim 4$ [Fig. 2(a)]. Consistently, we find that the Haar-averaged second SRE $\overline{M_{2}}$ is nearly indistinguishable from $\log(1-\overline{Y^{\rm lin}})$, confirming that the average in Eq. (2) is typical with vanishing fluctuations [49]. Thus we have
$\overline{M_{2}}=|A|+E+2+O(2^{-|A|-E}).$ (3)
Notice that $\overline{M_{2}}$ is extensive, as expected. Eqs. (2) and (3) establish a quantitative relation between the efficiency of local magic injection and the amount of stabilizer entanglement contained in the target state. For a fixed subregion $A$, $\overline{Y^{\rm lin}}(\rho)$ approaches $1$ exponentially with the amount of entanglement $E$ between $A$ and its complement, and correspondingly the average second SRE increases linearly with $E$, indicating that global magic injection on $AB$ is enhanced by preexisting entanglement. Similar behavior arises when $U_{A}$ is implemented by a brickwork circuit rather than a Haar-random unitary, although in that case the total injected magic is limited by the circuit depth (see SM [49]).
Intuitively, the enhancement of magic injection by stabilizer entanglement as revealed by Eqs. (2) and (3) can be understood as a consequence of gate ‘teleportation’ mediated by Bell pairs spanning subsystems $A$ and $B$ [see Fig. 1(c)]. As a simple example, consider a $2N$-qubit maximally entangled state between $A$ and $B$: $|{\rm Bell}\rangle_{AB}^{\otimes N}$. Then any unitary acting on half of the Bell pairs in $A$ will teleport to its partner in $B$ since $U\otimes I\ |\mathrm{Bell}\rangle_{AB}^{\otimes N}=I\otimes U^{T}\ |\mathrm{Bell}\rangle_{AB}^{\otimes N}$. It has been shown that any bipartite stabilizer state $|\psi\rangle_{AB}$ can be converted into a collection of single-qubit states (say $|+\rangle$) and Bell pairs through local Clifford unitaries [50]:
$U_{A}^{\rm Cliff}\otimes U^{\rm Cliff}_{B}|\psi\rangle_{AB}=|+\rangle^{\otimes f_{A}}|\mathrm{Bell}\rangle^{\otimes E}_{AB}|+\rangle^{\otimes f_{B}}.$ (4)
Hence for a stabilizer state that has entanglement, one can inject magic on $B$ through $A$ where magic flows to $B$ via the “entanglement conduit” formed by the $E$ Bell pairs. This simple picture also offers an intuitive explanation for the observed dynamics of local magic as a system thermalizes [26]—while time evolution increases local magic initially when the spins are unentangled, this local magic quickly spreads through the conduit as entanglement also grows in the system.
Figure 2: Numerical simulations of $\overline{Y^{\rm lin}}$ for the setup depicted in (a) Fig. 1(a), and (b) Fig. 1(b), respectively. We consider stabilizer initial states with varying system sizes and bipartite entanglement $E$. The analytical results Eqs. (2) and (LABEL:eq:UaUb) are shown as dashed lines.
@@
A further manifestation of entanglement-enhanced magic injection is revealed by applying a unitary that is a tensor product of local unitaries acting on subsytems $A$ and $B$: $U=U_{A}\otimes U_{B}$, as illustrated in Fig. 1(b). We compute the linear stabilizer entropy of the resulting state, averaged over Haar random $U_{A}$ and $U_{B}$. In the limit where $|A|\gg 1$ and $|B|\gg 1$, we find [49]
$\overline{Y^{\rm lin}}$ $=$ $\mathbb{E}_{U_{A}}\mathbb{E}_{U_{B}}Y^{\rm lin}\left(U_{A}U_{B}|\psi\rangle\langle\psi|U_{A}^{\dagger}U_{B}^{\dagger}\right)$
$=$ $1-4\cdot{2}^{-N}\left[1+3\cdot 2^{-2E}+O({{2}^{-|A|-E}},{{2}^{-|B|-E}})\right],$